# DeGauss: Dynamic-Static Decomposition with Gaussian Splatting for Distractor-free 3D Reconstruction

### v1

@@ Abstract
Reconstructing clean, distractor-free 3D scenes from real-world captures remains a significant challenge, particularly in highly dynamic and cluttered settings such as egocentric videos. To tackle this problem, we introduce DeGauss, a simple and robust self-supervised framework for dynamic scene reconstruction based on a decoupled dynamic-static Gaussian Splatting design. DeGauss models dynamic elements with foreground Gaussians and static content with background Gaussians, using a probabilistic mask to coordinate their composition and enable independent yet complementary optimization. DeGauss generalizes robustly across a wide range of real-world scenarios, from casual image collections to long, dynamic egocentric videos, without relying on complex heuristics or extensive supervision. Experiments on benchmarks including NeRF-on-the-go, ADT, AEA, Hot3D, and EPIC-Fields demonstrate that DeGauss consistently outperforms existing methods, establishing a strong baseline for generalizable, distractor-free 3D reconstruction in highly dynamic, interaction-rich environments. [^1] https://batfacewayne.github.io/DeGauss.io/
Figure 1: Given egocentric video frames or image collections with intense activities or transient objects, DeGauss yield state of the art results in modeling distractor-free static scene (a). Moreover, we demonstrate that DeGauss achieves high-quality dynamic scene modeling through a disentangled dynamic-static representation with consistently better mean LPIPS scores on test views (b).
@@ 3.6 Unsupervised scene decomposition
For the color mixing method in [15, 34, 18], the background and foreground renders are coupled during training, so updates to one branch constrain the expressiveness of the corresponding region in the other branch. This leads to local minima in case of initial misclassification of dynamic regions.
With our design, $\mathbf{C}_{b}$ update is controlled by $\mathbf{P}_{b}$ and independent from $\mathbf{C}_{f}$, which keeps the strong expressiveness of both branches while enabling clean distractor-free scene modeling with fine details.
@@ 4.3 Results
To assess the performance of our method for the distractor-free scene reconstruction task in the presence of noisy inputs, we conduct evaluations on both egocentric videos and image collections. For egocentric video data [17, 2, 21, 35]—which lack clean view references—we present qualitative comparisons with baseline methods [12, 25, 34] in Fig. 4. Compared to baseline methods [12, 34, 25], our method models high-quality distractor-free static background with accurate foreground separation. Moreover, we quantitatively compare the novel view synthesis quality between the composed dynamic-static render of NeuralDiff [34] and our methods on the test set. We additionally report video comparisons in our supplementary materials. For image collections dataset Nerf-on-the-go[23] with clean reference test views, we report detailed per-scene metrics including peak signal-to-noise ratio (PSNR), perceptual quality (LPIPS) [47], and structural similarity index (SSIM) [39] against baseline methods[24, 23, 12, 14, 25, 38] on the hold-out test set in Tab. 1. Our methods generalize to image collections and achieve state-of-the-art results. Notably, our method consistently achieves significantly better LPIPS scores over the previous SOTA method SpotlessSplats [25]. We show our method robustly handles occlusion and reconstructs fine static details compared to SpotlessSplats [25]in Fig. 5. Additionally, our methods could naturally handle various input challenges, such as camera motion blur and lens flare, as shown in Fig. 6.
@@ 5 Ablation study
Background Mask Element is introduced to promote cleaner separation and discourage mid-range probabilities. Though the improvements are not significant for image collections with good initializations, it leads to better dynamic-static modeling and separation results as shown in Fig. 9.
@@ Appendix D Implementation Details
Appendix E Additional Visualizations

### v2

@@ Abstract
Reconstructing clean, distractor-free 3D scenes from real-world captures remains a significant challenge, particularly in highly dynamic and cluttered settings such as egocentric videos. To tackle this problem, we introduce DeGauss, a simple and robust self-supervised framework for dynamic scene reconstruction based on a decoupled dynamic-static Gaussian Splatting design. DeGauss models dynamic elements with foreground Gaussians and static content with background Gaussians, using a probabilistic mask to coordinate their composition and enable independent yet complementary optimization. DeGauss generalizes robustly across a wide range of real-world scenarios, from casual image collections to long, dynamic egocentric videos, without relying on complex heuristics or extensive supervision. Experiments on benchmarks including NeRF-on-the-go, ADT, AEA, Hot3D, and EPIC-Fields demonstrate that DeGauss consistently outperforms existing methods, establishing a strong baseline for generalizable, distractor-free 3D reconstruction in highly dynamic, interaction-rich environments.
Figure 1: With self-supervised foreground-background gaussian splats modeling and accurate decomposition, DeGauss simultaneously enables (a): SOTA distractor-free static scene reconstruction for casual captures (no dynamic modeling in the static background) and (b): efficient, high-quality dynamic-static representation for dynamic scenes (no static modeling in dynamic foreground).
@@ 3.6 Unsupervised scene decomposition
Compositional rendering with color mixing in NeRF-based methods [16, 36, 19] sorts and integrates static and dynamic density and radiance along each ray(compose during rendering), leading to early ray termination during training on local minima and reconstructs static scene without fine details[36, 26]. In our decoupled design, the dynamic/static gaussians rasterize the foreground/background renders $\mathbf{C}_{f}$ and $\mathbf{C}_{b}$ independently and compose (after rending) with the probabilistic mask $\mathbf{P}_{f}$. This design enables full gradient flow and allow gradually formulated composition mask during training, as shown in Fig. 3. The wrongly modeled elements are gradually pruned during gaussian density control process, yielding accurate, clean yet flexible dynamic-static separation results that is much more robust to local minima.
@@ 4.3 Results
To assess the performance of our method for the distractor-free scene reconstruction task in the presence of noisy inputs, we conduct evaluations on both egocentric videos and image collections. For egocentric video data [18, 2, 23, 37]—which lack clean view references—we present qualitative comparisons with baseline methods [13, 27, 36] in Fig. 4. Compared to baseline methods [13, 36, 27], our method models high-quality distractor-free static background with accurate foreground separation. We additionally report video comparisons in our supplementary materials. For image collections dataset Nerf-on-the-go[25] with clean reference test views, we report detailed per-scene metrics including peak signal-to-noise ratio (PSNR), perceptual quality (LPIPS) [49], and structural similarity index (SSIM) [41] against baseline methods[26, 25, 13, 15, 27, 40] on the hold-out test set in Tab. 1. Our methods generalize to image collections and achieve state-of-the-art results. Notably, our method consistently achieves significantly better LPIPS scores over the previous SOTA method SpotlessSplats [27]. We show our method robustly handles occlusion and reconstructs fine static details compared to SpotlessSplats [27]in Fig. 5. Additionally, our methods could naturally handle various input challenges, such as camera motion blur and lens flare, as shown in Fig. 6.
@@ 5 Ablation study
Background Mask Element($m_{b}^{\prime}$) is introduced to promote cleaner separation and discourage mid-range probabilities. Though the improvements are not significant for image collections with good initializations, it leads to better dynamic-static modeling and separation results as shown in Fig. 9.
@@ 6 Conclusion
Acknowledgements. The authors would like to sincerely thank Siwei Zhang for the rigorous proof-reading and Hui Zhang for the discussion.
@@ Appendix D Implementation Details
Appendix E Efficiency Analysis on Neu3D [16] dataset
Our dynamic-static hybrid representation enables:(a) Much higher FPS: The time critical process of deformation prediction in 4DGS scales with the number of dynamic Gaussians. Table 4 shows we render 3× faster than 4DGS with superior quality by minimal dynamic element modeling with our dynamic-static decoupling design. (b) Better Quality We achieve much higher LPIPS and finer details in static (no stray motion) and dynamic (better handling disappearing gaussians), as reported in our paper and project page: https://batfacewayne.github.io/DeGauss.io/. Even on Coffee Martini, Flame Salmon with very far objects that poses challenges to gaussian splatting methods, our LPIPS and details remain best. (c) Applications The decoupled static with 3DGS seamlessly enables diverse applications as editing/styling.
Table 4: Quality and efficiency evaluation on all scenes of Neu3D [16] dataset tested on a RTX4090$\dagger$: trained densify grad threshold ×$\times 2$ to reduce number of gaussians.
Method PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) Training Time($\downarrow$) FPS($\uparrow$) Dyna. Gaussian num($\downarrow$)
NeRFPlayer [31] 30.29 0.909 0.151 6 hours 0.045 -
HyperReel [1] 30.72 0.931 0.101 9 hours 2.0 -
HexPlane [3] 30.00 0.922 0.113 12 hours 0.2 -
KPlanes [7] 31.63 0.964 0.117 5.0 hours 0.3 -
SWinGS [30] 31.12 0.941 0.095 - 71 -
4DGS [42] 31.12 0.937 0.058 0.85 hours 53 124,197
4DGS [42] 28.72 0.919 0.078 0.67 hours 68 62298
Ours 31.52 0.942 0.047 2.1 hours 71 56,533
Ours 31.56 0.942 0.047 2 hours 157 22,479
Appendix F Strict Monocular Input
Monocular reconstruction is extremely challenging. And compared to NeRF methods [24], dynamic gaussian methods [17, 42, 46] are highly expressive but much harder to regularize, generalizing poorly to novel views [32]Fig. 14). This actually enables our gaussian-based decoupled design, to fast and robustly separate dynamic/static modeling for a wide range of inputs, and our explicit static modeling leads to much better generalizability of novel view synthesis for dynamic scene modeling( Fig. 14).
Figure 14: Comparison with baseline methods on novel view synthesis with causal strict monocular input of dycheck-iphone dataset [8].
Appendix G Additional Experiment on Bonn RGBD dataset [22]
To further demonstrate generalizability our method, we evaluate on the Crowd scene of a SLAM dataset-Bonn RGBD [22], preprocessed with SFM and MVS pipeline of [28, 29]. We qualitatively show the distractor-free static scene reconstruction and dynamic-static decoupling results in Fig. 15.
Figure 15: Evaluation on 928 frames long highly dynamic Crowd scene of Bonn RGBD dataset [22](with only RGB as input).
Appendix H Additional Experiment on RobustNerf [26] dataset
We additionally report the performance of our method on RobustNerf [26] in Tab. 5 and Fig. 16.
Figure 16: Qualitative result on RobustNerf dataset [26].
Table 5: Quantitative results on RobustNerf [26] dataset. Our method shows best overall performance and significantly better LPIPS score over all baseline methods.
Method Android Crab2 Statue Yoda
[…]
3DGS [13] 23.32 0.794 0.159 31.76 0.925 0.172 20.83 0.830 0.148 28.92 0.905 0.192
WildGaussians [15] 24.67 0.828 0.151 30.52 0.909 0.213 22.54 0.863 0.129 30.55 0.905 0.202
SpotLessSplat [27] 24.20 0.810 0.159 33.90 0.933 0.169 21.97 0.821 0.163 34.24 0.938 0.156
Ours 24.54 0.813 0.083 34.48 0.952 0.076 23.08 0.861 0.097 33.48 0.947 0.082
Appendix I Additional Visualizations